If the atomic weight of carbon were set at 24 amu, the value of the Avogadro constant
would be?
1
Expert's answer
2013-01-24T07:39:18-0500
Avogadro constant enters into the definition of theunified atomic mass unit, u,
u = Mu / Na = 1,6605 * 10^-24
Mu- molar mass constant and Na is Avogadro constant.

Mu= Mw / Ar

Mu= 24 amu / 12 = 2 (If Mw = 24 amu)

Na'= Mu / 1,6605 * 10^-24 = 2 / 1,6605 * 10^-24 = 1,2
* 10^24

Thisis new Avogadro constant = 1,2 * 10^24